1) a) If a sample of pure hydrogen gas is ignited very carefully, the hydrogen burns gently, combining with the oxygen gas of the air to form water vapor. Write the unbalanced chemical equation for this reaction. (Include states-of-matter under the given conditions in your answer.)

b)Elemental boron (as a gas) is produced in one industrial process by heating diboron trioxide with magnesium metal, also producing magnesium oxide as a by-product. Write the unbalanced chemical equation for this process. (Omit states-of-matter from your answer.)
Answer #1

a) There is an unbalanced reaction:

H2 (g) + O2 (g) = H2O (g)

b) You have the reaction:

B2O3 + Mg = B + MgO

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.
Answer #2

Let's break down each part of the problem:

1a) Unbalanced Chemical Equation for Hydrogen Combustion

  • Reactants: Hydrogen gas (H₂) and Oxygen gas (O₂)

  • Product: Water vapor (H₂O)

The unbalanced equation is:

H₂(g) + O₂(g) → H₂O(g)

1b) Unbalanced Chemical Equation for Boron Production

  • Reactants: Diboron trioxide (B₂O₃) and Magnesium metal (Mg)

  • Products: Elemental Boron (B) and Magnesium oxide (MgO)

The unbalanced equation is:

B₂O₃ + Mg → B + MgO


source: self
answered by: anonymous
Answer #3

1). a)  :-

When pure hydrogen gas (H₂) is carefully ignited, it reacts with oxygen gas (O₂) from the air to form water vapor (H₂O). 

The unbalanced chemical equation is:


                           H₂(g) + O₂(g) → H₂O(g)

 



1). b) :-

In this industrial process, diboron trioxide (B₂O₃) is heated with magnesium metal (Mg) to produce elemental boron gas (B) and magnesium oxide (MgO) as a by-product. 

The unbalanced chemical equation is:


                          B₂O₃ + Mg → B + MgO


source: Self
answered by: Harshwardhan kunal